The following information from the national income accounts for a hypothetical country:
-------------------------------------------------------------------------------------------
GDP - $6000
Gross investment 800
Net investment 200
Consumption 4000
Government purchase of goods and services 1100
Government budget surplus 30
-------------------------------------------------------------------------------------------
What is
1)NDP
2)Net Exports
3)Government taxes minus transfers
4)Disposable personal income
5)Personal saving
1
Expert's answer
2014-12-17T11:38:58-0500
National income can be calculated by adding up all public and private expenditures made on final goods and services during a year. It is obtained by:
- Personal consumption expenditure of goods and services.
- Gross domestic private investment.
- Government purchase of goods and services.
- Net Foreign investment.
So we can measure national income using the following equation:
GDP = C + I + G + NE
So, we can find Net Exports as
NE = GDP - C - I - G
NE = 6000 – 4000 – 800 – 1100 = 100
NDP can be calculated:
NDP = GDP- Depreciation
Depreciation = Gross investment - Net investment = 800-200 = 600 NDP = 6000-600 = 5400 The next step is to calculate the government taxes minus transfers Government budget surplus = Government taxes minus transfers - Government purchase of goods and services
30 = Government taxes minus transfers -1100 Government taxes minus transfers = 30+1100 = 1130 Disposable personal income can be calculated:
Disposable Personal Income = National Income + Government Transfers to Individuals − Personal Tax - Depreciation= National Income – (Personal Tax - Government Transfers to Individuals) - Depreciation = 6000 -1130 – 600 = 4270 The next step is to calculate Personal saving:
Personal saving = Disposable Personal Income – Consumption = 270